OpenStudy (anonymous):

x^2 + 4=104 what is x? There will be a positive and negative number.

OpenStudy (john_es):

\[x^2=104-4=100\Rightarrow x^2=100\Rightarrow x=\pm\sqrt{100}=\pm10\]

OpenStudy (anonymous):

I am so confused

Directrix (directrix):

x^2 + 4=104 Subtract 4 from both sides. x^2 = 100 What number times itself is equal to 100? @Kaila_111697

Directrix (directrix):

Post what you think it is.

OpenStudy (anonymous):

10

Directrix (directrix):

If I had asked what number times itself equals 25, you would have said 5 because 5 times 5 equals 25. Then, I would have said but there is also a negative number that you can multiply times itself and get 25. You would have said -5 because -5 times -5 = 25.

Directrix (directrix):

@Kaila_111697 That is correct. What about the negative number.

OpenStudy (anonymous):

But it is not 100 its 104

OpenStudy (anonymous):

x^2+4=104 x^2=104-4=100 x=sqrt100=\[\pm10\]

Directrix (directrix):

@Kaila_111697 Look at the equation: x^2 + 4=104 We subtracted the 4 from both sides to get x^2 = 100.

OpenStudy (anonymous):

oh so 10 and -10

Directrix (directrix):

But 10 satisfies the equation: x^2 + 4=104 because 10*10 + 4 = 100 +4.

Directrix (directrix):

>>oh so 10 and -10 Yes.

OpenStudy (anonymous):

Thank you so much

Directrix (directrix):

Glad to help.